OpenStudy (xxgaming_mastaxx):

A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies, how many did they sell each week? How many boxes the first week? How many boxes the second week? How many boxes the third week

OpenStudy (igreen):

How many boxes the first week? 104 How many boxes the second week? 109 How many boxes the third week? 218

OpenStudy (igreen):

Let the original boxes sold be x so the no. of boxes sold the next week will be (x+5) and the boxes sold the 3rd week=2(x+5) total sale=431 boxes (given) x+(x+5)+2(x+5)=431 4x+15=431 4x=416 x=104

OpenStudy (xxgaming_mastaxx):

Thanks Man!